OpenStudy (abmon98):

express in partial fraction 2a/a^2-x^2

OpenStudy (anonymous):

this one is cooked up to be nice

OpenStudy (anonymous):

\[\frac{2a}{(a-x)(a+x)}=\frac{1}{a+x}+\frac{1}{a-x}\] you can check by adding

OpenStudy (mathstudent55):